OpenStudy (anonymous):

Solve for x: 2/5 (x − 4) = 2x.

OpenStudy (anonymous):

distribute the 2/5 to the stuff in parinthesese, so 2/5x-8/5=2x then subtract 2/5s from both sides 2/5x-8/5=2x -2/5x -2/5x so you get -8/5=-8/5x then multiply both sides by 5 to get -8=-8x then divide by negative 8 and get x=1 i would go back and plug it in to see if its right